### v1

# Neural Concept Verifier: Scaling Prover-Verifier Games via Concept Encodings

Abstract
While Prover-Verifier Games (PVGs) offer a promising path toward verifiability in nonlinear classification models, they have not yet been applied to complex inputs such as high-dimensional images. Conversely, Concept Bottleneck Models (CBMs) effectively translate such data into interpretable concepts but are limited by their reliance on low-capacity linear predictors. In this work, we introduce the Neural Concept Verifier (NCV), a unified framework combining PVGs with concept encodings for interpretable, nonlinear classification in high-dimensional settings. NCV achieves this by utilizing recent minimally supervised concept discovery models to extract structured concept encodings from raw inputs. A prover then selects a subset of these encodings, which a verifier—implemented as a nonlinear predictor—uses exclusively for decision-making. Our evaluations show that NCV outperforms CBM and pixel-based PVG classifier baselines on high-dimensional, logically complex datasets and also helps mitigate shortcut behavior. Overall, we demonstrate NCV as a promising step toward performative, verifiable AI.
Machine Learning, ICML
1 Introduction
Deep learning achieves remarkable predictive performance at the expense of interpretability and trustworthiness (Rudin,, 2019). However, particularly in high-stakes applications, it is critical that models provide verifiable justifications for their decisions (Irving et al.,, 2018; Fok and Weld,, 2023). Prover-Verifier Games (PVGs), introduced by Anil et al., (2021), formalize such justifications via a game-theoretic approach: a prover provides evidence to convince a verifier, who accepts only verifiable proofs. A prominent instantiation of PVGs is the Merlin-Arthur Classifier (Wäldchen et al.,, 2024), i.e., a classifier guided by cooperative and adversarial provers, offering formal interpretability guarantees through information-theoretic bounds. However, the Merlin-Arthur framework faces significant scalability challenges when applied to high-dimensional real-world data, as explanations based on raw pixels are both computationally difficult to optimize and offer limited human understandability (cf. Figure  1; Wäldchen et al., (2024)).
Figure 1: Challenges of Prover-Verifier Games (PVGs) in image classification: (i) It is non-trivial to scale up for high-dimensional data. (ii) Furthermore, the learned explanation masks on the pixel level remain difficult for humans to understand.
Figure 2: Overview of the Neural Concept Verifier (NCV). The input image is first processed by a concept extractor to produce symbolic concept encodings. A Prover–Verifier Game is then played over these encodings: a cooperative prover selects a sparse concept subset supporting the true class, while an adversarial prover selects misleading concepts. Finally, the verifier makes a prediction based only on these selected concepts, ensuring faithful and robust classification.
Concurrently, Concept Bottleneck Models (CBMs) have emerged as a powerful framework for interpretable machine learning, structuring predictions through intermediate interpretable concept encodings (Koh et al.,, 2020; Stammer et al.,, 2021). Despite their advantages, CBMs typically employ linear classifiers on top of concept encoding layers, thereby potentially restricting expressivity and failing on tasks requiring nonlinear interactions among concepts (e.g., XOR problems, counting or permutation invariance) (Mahinpei and et al.,, 2021; Kimura et al.,, 2024; Lee et al., 2019b, ).
In this work, we bridge these two complementary directions by introducing the Neural Concept Verifier (NCV), a novel framework integrating concept-based representations into PVGs in the form of Merlin-Arthur Classifiers. NCV shifts the prover–verifier interaction from the image level to a structured, symbolic concept level, overcoming both the scalability limitations encountered by Merlin-Arthur Classifiers in high-dimensional settings and the expressivity constraints inherent to linear Concept Bottleneck Models. Through extensive evaluations on controlled synthetic benchmarks such as CLEVR-Hans, as well as large-scale real-world datasets including CIFAR-100 and ImageNet-1k, we demonstrate that NCV successfully scales PVGs to complex, high-dimensional classification tasks. Additionally, NCV enables verifiable, performant nonlinear classifiers on top of concept extractors, effectively narrowing the interpretability–accuracy gap present in standard, linear CBMs. Lastly, our framework improves robustness to shortcut learning, thereby enhancing the generalizability and trustworthiness of predictions, particularly crucial in high-stakes applications.
Our contributions can be summarised as: (i) We propose Neural Concept Verifier (NCV), a framework combining concept-based models with Prover–Verifier Games (PVGs). (ii) NCV scales PVGs to high-dimensional image data by operating on compact concept encodings. (iii) It enables expressive yet interpretable classification via sparse, nonlinear reasoning over concepts. (iv) We validate NCV on synthetic and real-world benchmarks, demonstrating strong accuracy and verifiability. (v) We highlight that NCV improves generalization under spurious correlations, indicating increased robustness.
The remainder of the paper is structured as follows. We begin with a review of related work, highlighting recent developments in the field. We then introduce our proposed Neural Concept Verifier framework, and present its formal description. This is followed by a comprehensive experimental evaluation that investigates key aspects of the framework. Finally, we discuss our findings and conclude the paper.
2 Background
Prover-Verifier Games. PVGs were introduced by Anil et al., (2021) as a game-theoretic framework to encourage learning agents to produce testable justifications through interactions between an untrusted prover and a trusted verifier. Their work showed that under suitable game settings, the verifier can learn robust decision rules even when the prover actively attempts to persuade it of arbitrary outputs. Wäldchen et al., (2024) extended this idea to the Merlin-Arthur Classifier (MAC), which provides formal interpretability guarantees by bounding the mutual information between the selected features and the ground-truth label. Recently, Kirchner et al., (2024) applied a PVG-inspired approach to improve legibility of Large Language Model (LLM) outputs and Amit and et al., (2024) introduced self-proving models that leverage interactive proofs to formally verify the correctness of model outputs. PVG-style setups have also been explored in safety-focused learning protocols (Irving et al.,, 2018; Brown-Cohen et al.,, 2024; Głuch et al.,, 2024). These developments reflect a broader trend of utilising multi-agent learning (Pruthi et al.,, 2022; Schneider and Vlachos,, 2023; Du et al.,, 2023; Nair et al.,, 2023; Stammer et al., 2024a, ). Our work builds on this line of research by embedding PVGs into a concept-based classification framework, addressing the dimensionality bottleneck that limits PVGs in high-dimensional settings.
Concept Representations for Interpretability. The introduction of Concept Bottleneck Models (CBMs) (Koh et al.,, 2020; Delfosse et al.,, 2024), but also concept-based explanations (Kim et al.,, 2018; Crabbé and van der Schaar,, 2022; Poeta et al.,, 2023; Lee et al.,, 2025) was an important moment in the growing interest in AI interpretability research. The particular appeal of CBMs lies in the promise of interpretable predictions and a controllable, structured interface for human interactions (Stammer et al.,, 2021). While initial CBMs relied on fully supervised concept annotations, subsequent research has relaxed this requirement by leveraging pretrained vision-language models like CLIP for concept extraction (Bhalla et al.,, 2024; Yang et al.,, 2023; Oikarinen et al.,, 2023; Panousis et al.,, 2024), or employing fully unsupervised concept discovery methods (Ghorbani et al.,, 2019; Stammer et al., 2024b, ; Schut et al.,, 2025; Sawada and Nakamura,, 2022). While previous work has focused on enhancing the concept bottleneck itself, by reducing supervision requirements, mitigating concept leakage (Havasi et al.,, 2022), or dynamically expanding the concept space (Shang et al.,, 2024), the classifier component of CBMs, has received little attention. We argue that while a linear layer offers interpretability, its limited expressiveness makes it unsuitable for handling complex, compositional, and object-centric tasks. Our proposed NCV tackles this limitation.
Shortcut Learning. Independent of their interpretability, training deep models can often lead to unwanted artifacts and side effects. Shortcut learning describes the problem of models learning to rely on unwanted and unintended features to resolve a task (Geirhos et al.,, 2020). This is a common problem when training purely deep learning models (Lapuschkin et al.,, 2019; Schramowski et al.,, 2020) or neuro-symbolic models (Marconato et al.,, 2023; Bortolotti et al.,, 2025), and, if not taken care of, can lead to predictions being right for the wrong reasons (Ross et al.,, 2017). There have been various methods developed to tackle this problem, from careful dataset curation (Ahmed et al.,, 2021) to modified model training (Friedrich et al.,, 2023), cf. (Steinmann et al.,, 2024) for a comprehensive overview. However, these mitigation methods rely on several important assumptions about the training data and potential shortcuts to mitigate, and can affect model performance (Sagawa et al.,, 2019). While NCV is not specifically designed to mitigate shortcuts, we show that our setup can intrinsically mitigate the impact of shortcuts in the data.
3 Neural Concept Verifier (NCV)
In this section, we introduce the Neural Concept Verifier (NCV), a novel framework for training interpretable and verifiable nonlinear classifiers capable of handling high-dimensional inputs. NCV combines concept-based representations with the Prover-Verifier Game setup of MAC (Wäldchen et al.,, 2024), enabling models whose predictions are certified to rely on a sparse set of high-level concepts. This approach supports both model interpretability and predictive performance, making concept-based models more competitive with opaque architectures and at the same time extending the benefits of PVGs to high-dimensional data. NCV consists of two main components: (i) a minimally or weakly supervised concept extractor that transforms input data into interpretable concept encodings, and (ii) MAC that selects and verifies sparse concept-based inputs to make final predictions. An overview of the framework is shown in Figure 2. After providing background notations and a high-level overview of the NCV framework below, we provide detailed descriptions of each main component as well as training and inference details in the following.
3.1 Problem Setup and Notation
Let $\mathcal{X}\in\mathbb{R}^{N\times D}$ denote a dataset of $N$ inputs (e.g., images), each of dimension $D$, and $\mathcal{Y}\in\mathbb{R}^{N\times(K+1)}$ the corresponding labels for $K$ classes plus one additional rejection class. The rejection class allows the model to abstain from a decision when uncertain, which is important for enforcing interpretability guarantees in adversarial setups (Wäldchen et al.,, 2024). Our overall goal is to learn a model $f:\mathcal{X}\rightarrow\mathcal{Y}$ that maps an input $x\in\mathcal{X}$ to a prediction $\hat{y}\in\mathcal{Y}$, where the prediction is based on a small, interpretable subset of high-level concepts.
3.2 The NCV Framework
NCV next decomposes $f$ into the following components (cf. Figure  2 for an illustrative overview):
  1. 1.
    A concept extractor $g:\mathcal{X}\rightarrow\mathcal{C}$, which maps each input $x$ to a high-level concept encoding $\mathbf{c}\in\mathbb{R}^{C}$, where $C\in\mathbb{N}$ is the number of discovered concepts.
  2. 2.
    A pair of prover agents $M,\widehat{M}:\mathcal{C}\rightarrow\{0,1\}^{C}$ that produce sparse binary masks selecting $m$ concepts each. $M$ (Merlin, cooperative Prover) aims to help classification; $\widehat{M}$ (Morgana, adverserial Prover) aims to mislead.
  3. 3.
    A verifier (Arthur) $A:\mathbb{R}^{C}\rightarrow\mathcal{Y}$, which predicts a label based only on the masked concepts.
The three agents are trained jointly, where the interaction between these agents encourages the verifier to rely only on robust, informative concept features. Let us now provide details on these individual components.
3.3 Concept Extraction
The concept extractor $g$ transforms input images into interpretable, high-level feature encodings. While NCV is compatible with any concept bottleneck or discovery model, we experiment with two instantiations in our evaluations:
CLIP-based Concept Extraction.
In one instantiation we leverage pretrained vision-language models such as CLIP (Radford et al.,, 2021). Given an image $x$ and a vocabulary of textual concept descriptions $T=\{t_{1},...,t_{C}\}$, we compute: $\mathbf{c}_{i}=\langle e_{\text{img}}(x),e_{\text{text}}(t_{i})\rangle\quad\forall i\in\{1,...,C\}$, where $e_{\text{img}}$ and $e_{\text{text}}$ are CLIP’s image and text encoders, respectively. The resulting concept encoding $\mathbf{c}\in\mathbb{R}^{C}$ measures the similarity of the image to each textual concept. Unlike sparse activation methods such as SpLiCE (Bhalla et al.,, 2024), which enforce sparsity directly in the concept activation space $\mathbf{c}$ to promote human interpretability, we shift this constraint to the level of concept usage. Thus, in NCV, we maintain a dense and expressive concept space but enforce sparsity at the downstream classifier level, where only a small subset of concepts is selected for each decision. This results in task-driven sparsity that emerges naturally through the prover–verifier interaction.
Unsupervised Concept Extraction via NCB.
With our second instantiation we demonstrate that NCV is also compatible with unsupervised concept discovery modules. For this, we instantiate $g$ using the Neural Concept Binder (NCB) (Stammer et al., 2024b, ) that infers discrete, symbolic concepts from raw images $\mathcal{X}$, without any supervision. It first produces soft, continuous object-centric encodings $z=g_{\text{soft}}(x)$ using slot-based attention and factor binding. These are then discretized via retrieval-based clustering, resulting in hard, symbolic concept encodings $\mathbf{c}=g_{\text{hard}}(z)$, leading to $\mathbf{c}=(g_{\text{hard}}(g_{\text{soft}}(x))\in\mathbb{R}^{C}$, where each entry corresponds to a learned visual concept. Importantly, NCB’s concept encodings (in comparison to those extracted via CLIP) are object-centric, i.e., represent concepts per object in an image. Since NCB does not rely on textual prompts or class labels, its concept space may include noisy or redundant concepts. However, the sparse concept selection in NCV mitigates this by identifying only the most relevant concepts for each prediction.
The goal of investigating these two concept extractor instantiations is to demonstrate NCV’s generality and provide a good testbed to investigate NCV in the context of real-world, natural images, but also object-centric concept settings.
3.4 Verifiable Classification via the Merlin-Arthur Classifiers
The second core component of NCV is a verifiable classifier of the Merlin-Arthur (MA) setup (Wäldchen et al.,, 2024) originally inspired by Interactive Proof Systems (Goldwasser et al.,, 1989). This setup generally formalizes the idea of proving that a classification decision is supported by a sparse and informative set of features.
In NCV, specifically, two competing provers, Merlin and Morgana, select concept subsets either to support or mislead classification, respectively. The verifier, Arthur, then makes predictions based solely on these masked concepts (often denoted as “certificates” in PVG literature) without knowledge of the selector’s intent. Formally, each prover outputs a sparse binary mask with $m$ active entries, producing selected subsets:
$\mathcal{S}=M(\mathbf{c})\odot\mathbf{c},\quad\widehat{\mathcal{S}}=\widehat{M}(\mathbf{c})\odot\mathbf{c},$
where $\mathbf{c}$ is the concept encoding from extractor $g$, and $\odot$ denotes element-wise masking. Notably, all three agents represent differentiable neural models, with Arthur specifically representing a non-linear model.
This interactive setup enables two key metrics: (i) completeness, i.e., the probability that Arthur correctly classifies the input when using Merlin’s selected features and (ii) soundness, i.e., the probability that Arthur avoids incorrect classification when using Morgana’s features, either by giving the correct class or opting for a rejection. Overall, these components in NCV ensure that only sparse, faithful concept subsets are used and can be inspected for final predictions.
3.5 Training and Inference
The training step of NCV incorporates updating only the parameters of the three agents $M$, $\widehat{M}$, and $A$ as $g$ represents a pretrained model that is subsequently frozen upon NCV’s multi-agent training step. Thus, the three agents are jointly trained by optimizing a three-agent game, which encourages Arthur to rely on concepts selected by Merlin, while being robust to potentially misleading concepts selected by Morgana. Given a concept encoding $\mathbf{c}\in\mathbb{R}^{C}$, label $y\in\mathcal{Y}$ and cross-entropy function $CE(\cdot,\cdot)$, we define:
  • Merlin’s loss:$L_{M}=CE(A(\mathcal{S}),y)$, where $\mathcal{S}$ is the sparse concept subset selected by Merlin. This loss encourages Arthur to classify correctly based on Merlin’s input.
  • Morgana’s loss:$L_{\widehat{M}}=CE(A(\widehat{\mathcal{S}}),y)$, where $\widehat{\mathcal{S}}$ is Morgana’s adversarial concept subset. Here, the loss is interpreted as the classifier’s inability to be misled by deceptive inputs [^1] In practice, the CE loss of Morgana is a slightly modified CE loss, cf. Wäldchen et al., (2024)..
Overall, Arthur’s loss combines both objectives with a hyperparameter $\gamma\in\mathbb{R}_{\geq 0}$, controlling the emphasis on predictive performance (completeness) versus robustness (soundness):
$L_{A}=(1-\gamma)\,L_{M}+\gamma\,L_{\widehat{M}},$ (1)
In detail, the three agents are updated jointly in a two-phase min-max optimization. First, the prover agents are updated where Merlin minimizes $L_{M}$ and Morgana maximizes $L_{\widehat{M}}$; then Arthur is updated by minimizing $L_{A}$ on the sparse selected concepts chosen by the provers. This scheme incentivizes Arthur to base its predictions on informative, task-relevant, and verifiably robust concept subsets.
At inference time, only the cooperative prover $M$ is used to select a sparse subset of concepts, based on the input’s concept encoding. The verifier $A$ then predicts a label or rejects based solely on this selected subset.
Overall, by integrating concept-extractor modules and leveraging the Merlin–Arthur framework, NCV emphasizes faithfulness and interpretability while preserving nonlinear modeling capabilities, and shifts the min–max optimization into a lower-dimensional concept space, improving efficiency, scalability, and stability in high-dimensional settings—a common challenge in min–max optimization for deep learning (Mescheder et al.,, 2018; Nagarajan and Kolter,, 2017).
4 Experimental Evaluations
In this section, we present a comprehensive evaluation of Neural Concept Verifier (NCV) on both synthetic and real-world high-dimensional image datasets. We evaluate based on two instantiations of NCV that utilize different concept extractors: a CLIP-based extractor and the Neural Concept Binder (NCB). We assess predictive performance and interpretability across multiple datasets, compare against established baselines, including SpLiCE and pixel-based Merlin-Arthur Classifiers, and examine scalability and robustness against shortcut learning.
Our evaluation is structured around the following research questions: (Q1) Does shifting Prover-Verifier Games (PVGs) to concept-encodings via NCV lead to performative classifiers on high-dimensional images (i.e., high completeness and soundness)? (Q2) Does NCV reduce the “interpretability-accuracy gap” in the context of CBMs? (Q3) Does NCV allow for more detailed explanations over pixel-based PVGs? Finally, (Q4) Can training via NCV reduce shortcut learning?
4.1 Experimental Setup
4.1.1 Datasets.
We investigate NCV in the context of the CLEVR-Hans3 and CLEVR-Hans7 (Stammer et al.,, 2021) datasets, which are synthetic benchmarks derived from CLEVR (Johnson et al.,, 2017) that capture complex object compositions and include visual shortcuts. CLEVR-Hans3 features three object classes, while CLEVR-Hans7 increases the task complexity to seven object classes. Both datasets consist of images rendered at a resolution of $128\!\times\!128$ pixels. The original datasets contain shortcuts from attributes to class labels in their training and validation sets, which are not present in the test set, making them well-suited for studying shortcut behavior. For example, in CLEVR-Hans3, a class may be defined by object shape and size, but during training, irrelevant features like color (e.g., gray cubes) are spuriously correlated with the label. Models relying on such shortcuts can fail under the decorrelated test distribution. We first evaluate our results on non-confounded versions of these datasets, where feature distributions are consistent across training, validation, and test splits. We return to the original confounded datasets when investigating shortcut mitigation.
Further, to assess scalability and generalization to natural image data, we evaluate on the widely-used benchmarks ImageNet-1k (Deng et al.,, 2009) and CIFAR-100 (Krizhevsky,, 2009). ImageNet-1k consists of over 1.2 million high-resolution images across 1,000 classes, typically resized to $224\!\times\!224$ pixels for training. CIFAR-100 provides 60,000 low-resolution images of $32\!\times\!32$ pixels across 100 fine-grained categories, representing a more compact but still challenging dataset to evaluate generalization.
4.1.2 Baselines
We compare our proposed framework against several representative baselines. Training parameters and implementation details for all baselines are provided in Suppl.  A.
Black-box Classifier: ResNet-18. We first consider a standard ResNet-18 model (He et al.,, 2016) trained end-to-end on raw image inputs (denoted as ResNet-18). This serves as a high-performance but non-interpretable baseline model.
Merlin-Arthur Classifier on Pixel-Space (Pixel-MAC). We further evaluate an instantiation of a Prover-Verifier Game, as a pixel-based Merlin-Arthur Classifier (Wäldchen et al.,, 2024) (denoted as Pixel-MAC). Hereby, the verifier is initialized from a pretrained ResNet-18 model, whereas both the cooperative prover (Merlin) and adversarial prover (Morgana) are implemented as U-Net models (Ronneberger et al.,, 2015) that output continuous-valued feature importance masks over the input image. These masks are subsequently discretized using a Top-$k$ selection to define the subset of features visible to the verifier. All agents are jointly fine-tuned within the prover-verifier setup. The explanations (i.e., certificates) in this setup represent masks on the original pixel space. For a detailed description of the general Pixel-MAC workflow, we refer to (Wäldchen et al.,, 2024).
Concept Bottleneck Model with Linear Classifier (CBM). Lastly, we compare NCV to a vanilla CBM setup in which the final classifier represents a linear model (denoted as CBM). In these baselines, a concept extractor first generates a vector of interpretable concept features from the input image. The resulting concept features are then passed to a linear classifier trained to predict the target class. We leverage NCB (Stammer et al., 2024b, ) as the concept extractor for evaluations on CLEVR-Hans and SpLiCE (Bhalla et al.,, 2024) as the concept extractor for evaluations on CIFAR-100 and ImageNet-1k.
4.1.3 NCV Model Instantiations
NCB-based Neural Concept Verifier. As detailed in Section  3.3, for evaluations on CLEVR-Hans3 and CLEVR-Hans7, we utilize NCB (Stammer et al., 2024b, ) to extract unsupervised, discrete concepts. Specifically, we use NCB models that were pretrained on the CLEVR (Johnson et al.,, 2017) dataset. NCB’s resulting concept encodings are structured as a set of object-centric slots. To handle this structure, we employ a permutation-invariant Set Transformer (Lee et al., 2019a, ) as the verifier (Arthur). The provers (Merlin and Morgana) are implemented as independent neural networks, each parameterized by a Set Transformer. They receive the full concept slot matrix as input and output a sparse selection mask with 12 nonzero entries, indicating the active concept blocks provided to the verifier. All components are jointly trained using the Adam optimizer. Further details and ablation studies can be found in Suppl.  B.
CLIP-based Neural Concept Verifier. For evaluations on ImageNet-1k and CIFAR-100, we employ an instantiation of NCV with a CLIP-based concept extractor (Radford et al.,, 2021). We follow the approach of SpLiCE (Bhalla et al.,, 2024), which computes image-to-text similarity scores between CLIP image features and a predefined vocabulary of candidate concepts. However, where SpLiCE performs a per-sample optimization to obtain a sparse activation of the concepts, our NCV instantiation retains the full concept activation vector (denoted as CLIP-Sim in the following) and delegates the selection of relevant concepts to the subsequent prover agent in our NCV framework. This avoids expensive optimization at inference time and allows our approach to scale efficiently while preserving the ability to select meaningful concepts for interpretable classification. The verifier, Merlin and Morgana are all implemented as two-layer multilayer perceptron (MLP). The two provers receive the full concept activation vector as input and output a sparse selection mask of $32$ concepts per example. All components are respectively trained using the Adam optimizer. Further details and ablation studies can be found in Suppl.  C.
4.1.4 Metrics.
All methods are evaluated in terms of completeness and soundness (where applicable), as described in Section  3.4. Note that completeness is equivalent to standard classification accuracy in our setting, as it evaluates the model’s accuracy under cooperative feature selection, whereas soundness describes the robustness against adversarial prover selections. We evaluate with 20 random seeds for the CLEVR-Hans evaluations and 10 seeds for ImageNet-1k and CIFAR-100. We report the average performance and standard deviation over these seeds in all tables. Furthermore, when evaluating shortcut learning in the context of the CLEVR-Hans datasets, we report the validation–test gap, defined as the difference between validation accuracy on a confounded split and test accuracy on a non-confounded split. A smaller gap indicates improved generalization and reduced reliance on spurious correlations.
Table 1: NCV delivers high predictive performance and robustness through verifiable, concept-based reasoning. We report completeness and soundness scores for ResNet-18, Pixel-MAC, CBM, and NCV across synthetic (CLEVR-Hans3, CLEVR-Hans7) and real-world (CIFAR-100, ImageNet-1k) datasets. NCV consistently matches or outperforms baselines in completeness while offering strong soundness guarantees.
Model Feature Space Completeness (Accuracy) Soundness (Robustness)
CLEVR-Hans3
[…]
CBM NCB 95.44± 0.0895.44plus-or-minus0.0895.$4$ 0.08 n/a
NCV (ours) NCB 98.92± $\boldsymbol{98.92\mbox{\scriptsize$\pm\,0.32$}}$ 100.00± $\boldsymbol{100.00\mbox{\scriptsize$\pm\,0.00$}}$
CLEVR-Hans7
[…]
NCV (ours) NCB 97.89± $\boldsymbol{97.89\mbox{\scriptsize$\pm\,0.31$}}$ 100.00± $\boldsymbol{100.00\mbox{\scriptsize$\pm\,0.00$}}$
CIFAR-100
ResNet-18 pixel space 79.73 ± $\pm\,0.36$ n/a
[…]
CBM SpLiCE 75.42 ± $\pm\,0.04$ n/a
NCV (ours) CLIP-Sim 83.32± $\boldsymbol{83.32\mbox{\scriptsize$\pm\,0.28$}}$ 99.99± $\boldsymbol{99.99\mbox{\scriptsize$\pm\,0.01$}}$
ImageNet
ResNet-18 pixel space 66.16± $\pm\,0.41$ n/a
[…]
CBM SpLiCE 68.59± $\boldsymbol{68.59\mbox{\scriptsize$\pm\,0.01$}}$ n/a
[…]
4.2 Evaluations
Scalability to High Dimensions (Q1).
In our first evaluation, we examine whether shifting the Prover–Verifier Game (PVG) to concept encodings enables NCV to scale to high-dimensional image domains while achieving strong performance in terms of completeness and soundness. We hereby compare NCV against two key baselines: (1) a black-box ResNet-18 classifier, and (2) Pixel-MAC, a nonlinear PVG model operating in raw pixel space. Table  1 summarizes results across synthetic (CLEVR-Hans3, CLEVR-Hans7) and real-world (CIFAR-100, ImageNet-1k) benchmarks. Each model’s feature space is indicated for clarity. On the synthetic CLEVR-Hans benchmarks, we observe that NCV consistently achieves the highest completeness scores, surpassing Pixel-MAC and even ResNet-18 on CLEVR-Hans3, while also attaining perfect soundness. This demonstrates that NCV not only matches or exceeds the performance of strong black-box classifiers but also maintains faithful and certifiable decision-making. Pixel-MAC performs well in these settings but falls slightly short in completeness and cannot match NCV’s zero-error soundness.
On the more challenging real-world datasets CIFAR-100 and ImageNet-1k, Pixel-MAC either fails entirely or performs poorly. In contrast, NCV successfully scales to these datasets, achieving superior completeness and near-perfect soundness without incurring test-time optimization overhead. Notably, NCV surpasses ResNet-18 even in raw accuracy, providing both higher predictive performance while retaining interpretability. In summary, NCV generalizes well across domains: it scales beyond the limitations of pixel-based PVGs, delivers competitive accuracy even on large-scale and complex datasets, and retains soundness throughout. These findings affirm that shifting PVGs to concept space enables interpretable and verifiable classifiers that remain performant and scalable in high-dimensional visual environments. We therefore answer Q1 affirmatively.
Narrowing the Interpretability–Accuracy Gap (Q2).
In Table  1, we further examine whether NCV can overcome a central limitation of standard Concept Bottleneck Models (CBMs): the interpretability–accuracy gap resulting from their use of constrained linear classifiers. Across all evaluated datasets, NCV consistently narrows—and in some cases closes—this gap, while maintaining formal verifiability. On CLEVR-Hans3, CBM trails the opaque ResNet-18 by over 2 percentage points in completeness, whereas NCV matches or exceeds the ResNet’s performance while retaining perfect soundness. The benefit is even more pronounced on CLEVR-Hans7: CBM underperforms ResNet-18 by nearly 10 percentage points, while NCV narrows the gap to just 1 percentage point and again delivers flawless soundness. This trend persists on real-world datasets. On CIFAR-100, NCV outperforms CBM by nearly 8 percentage points in completeness and even slightly exceeds ResNet-18’s performance, while providing robust and interpretable predictions. On ImageNet-1k, CBM slightly outperforms NCV in raw completeness by a small margin. Overall, NCV consistently improves over linear CBMs in both accuracy and robustness, especially on tasks requiring complex concept reasoning. At the same time, it matches or even surpasses the accuracy of the ResNet-18 model, demonstrating that interpretable, concept-level reasoning via Prover–Verifier Games can deliver competitive performance without sacrificing transparency or verifiability. We therefore answer Q2 affirmatively.
Higher Detailed Explanations (Q3)
We next investigate the resulting explanations produced by our NCV framework, with a focus on interpretability. Since our goal is to improve the explanatory clarity of Prover–Verifier Games, we compare against pixel-level MAC explanations, which serve as a natural baseline for interpretability within the PVG framework. Figure  3 illustrates a qualitative example from the CLEVR-Hans3 dataset (cf. Figure  4 for more examples), both from the Pixel-MAC setup and NCV. Notably, under Pixel-MAC (see top panel of Figure  3) the Prover–Verifier setup operates directly on pixels, yielding broad, diffuse masks that often cover entire objects or irrelevant background regions, arguably providing limited insight regarding which exact features drive the verifier’s final decision. These patterns also reflect typical failure modes of Pixel-MAC despite high completeness and soundness. In contrast, NCV leverages its internal concept encodings $\mathbf{c}$ to pinpoint the precise attributes relevant for a class decision (see bottom panel of Figure  3), e.g., a “small metal cube” and “small metal sphere”. [^2] It is important to note that this object-level specification is dependent on the utilized concept extractor. NCB, by design allows to receive object-specific concept explanations. However, this is by default not possible via the CLIP-based concept extractor. In any case, NCV allows for receiving explanations on the level of its concept-encodings. Overall, these examples highlight that NCV offers substantially more detailed and understandable explanations than pixel-based PVG baselines, highlighting that concept-level PVGs yield verifiable decisions that are potentially aligned with underlying logical rules. This leads us to answer Q3 affirmatively.
Figure 3: Comparison of explanations from NCV vs. Pixel-MAC for a CLEVR-Hans3 image (class label two) with GT class rule: small metal cube and small metal sphere. (a) Merlin–Arthur training on pixel space yields uninformative masks. (b) MAC on concept encodings via NCV provides clear explanations by highlighting object features corresponding to the class rule. The single-object images are reconstructions from the respective slots selected by Merlin (prover).
Mitigating Shortcut Learning (Q4)
Lastly, to assess whether NCV can mitigate shortcut learning in image classification, we train models on different versions of CLEVR-Hans7 with varying ratios of clean samples (i.e., without shortcut) in the training and validation sets. We then measure validation accuracy with shortcuts and test accuracy on a held-out, clean data split. This setup allows us to track both predictive performance and robustness to shortcut learning. Table  2 reports results for three model types: a linear CBM, a non-linear CBM, and our instantiation of NCV using NCB as concept extractor. We observe that while NCV achieves the highest test accuracy among all models in the 0% clean data setting, it still exhibits a sizeable validation-test gap, indicating a strong influence of the underlying shortcuts. Notably, NCV shows a smaller gap than the non-linear CBM, suggesting that the inductive bias of the PVG-based training objective encourages a smaller influence of the shortcut, even without access to clean training data. As we progressively increase the amount of clean samples (from 1% to 20%), test accuracy and test-validation gap improves across all models. However, NCV consistently achieves the highest test accuracy in every setting, and its validation–test gap decreases more rapidly than for either CBM variant. This trend indicates that NCV is not only better at leveraging clean supervision when available, but is also more robust to shortcut learning. A similar trend is observed for CLEVR-Hans3 (cf. Table  6), where NCV closes the generalization gap even more efficiently than all baselines. Importantly, the linear CBM approach maintains a much larger gap than both non-linear approaches, with NCV providing additional soundness guarantees over the non-linear CBM. Together, these results demonstrate that concept-level Prover–Verifier Games in NCV encourage models to base decisions on more robust, task-relevant features. As a result, NCV is more resilient to shortcut learning, inparticular with access to minimal amounts of clean data.
Table 2: NCV shows increased robustness to shortcuts on CLEVR-Hans7. We report validation accuracy on a set with shortcuts, and test accuracy on a set without shortcuts, as well as the validation–test gap (lower is better) across models trained with varying amounts of clean data.
Ratio Clean (Samples) Model Val Acc (w/ shortcut) Test Acc (w/o shortcut) Val-Test Gap (↓)
[…]
NCV 98.38± $\mathbf{98.38}\mbox{\scriptsize$\pm\,0.18$}$ 92.23± $\mathbf{92.23}\mbox{\scriptsize$\pm\,0.67$}$ $6.15$
[…]
NCV 98.19± $\mathbf{98.19}\mbox{\scriptsize$\pm\,0.24$}$ 94.68± $\mathbf{94.68}\mbox{\scriptsize$\pm\,0.64$}$ $\mathbf{3.51}$
[…]
NCV 98.47± $\mathbf{98.47}\mbox{\scriptsize$\pm\,0.24$}$ 96.24± $\mathbf{96.24}\mbox{\scriptsize$\pm\,0.71$}$ $\mathbf{2.23}$
[…]
NCV 98.63± $\mathbf{98.63}\mbox{\scriptsize$\pm\,0.13$}$ 97.74± $\mathbf{97.74}\mbox{\scriptsize$\pm\,0.28$}$ $\mathbf{0.89}$
5 Discussion
Overall, our results show that shifting Prover–Verifier Games (PVGs) to the concept level yields a powerful and scalable framework for verifiable, interpretable classification. By operating on symbolic concept embeddings, NCV avoids the computational cost of per-sample inference in pixel space, yet matches or surpasses pixel-based baselines in both completeness and soundness. It reduces the performance gap typical of Concept Bottleneck Models (CBMs), achieving parity with opaque models on synthetic tasks and even surpassing them on natural images. Additionally, concept-level outputs offer concise, human-readable explanations. Finally, NCV exhibits a resilience to spurious correlations, generalizing from confounded training splits and closing the generalization gap with minimal injections of clean data.
That said, NCV has several limitations. Its effectiveness depends on the quality of the underlying concept extractor: noisy or entangled concept vocabularies can reduce both accuracy and human understandability. The increased training complexity introduced by the three-agent PVG setup also results in greater computational cost and training instability, e.g., compared to linear CBMs. Moreover, when using pretrained models like CLIP for concept discovery, NCV inherits their biases and inconsistencies to some extent (Birhane et al.,, 2021; Gehman et al.,, 2020; Bhalla et al.,, 2024). Finally, recent work (Debole et al.,, 2025) shows that such concept spaces can diverge from expert semantics, even when yielding strong downstream performance.
6 Conclusion
In this work, we have introduced the Neural Concept Verifier (NCV), a unified framework that brings together Prover–Verifier Games and concept-level representations for interpretable classification at scale. Through extensive experiments on CLEVR-Hans, CIFAR-100, and ImageNet-1k, we have shown that NCV achieves high completeness and soundness, reduces the interpretability–accuracy gap of concept bottleneck models, delivers detailed concept-based explanations, and effectively mitigates shortcut learning. Thus, NCV paves the way for deploying trustworthy and transparent models in domains where both predictive performance and verifiability are essential.
Future work should explore how concept encodings can be integrated into alternative PVG-style setups, where structured representations may improve performance or reduce communication overhead. It is also promising to investigate applications beyond vision, such as natural language processing and structured data, where interpretable verification may be equally valuable. Finally, while existing information-theoretic guarantees, such as those introduced by Wäldchen et al., (2024), focus on binary classification under specific assumptions, extending such guarantees to high-dimensional, multi-class settings remains an important open direction for formal interpretability at scale.
Acknowledgements
This research was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689). This work was further supported by the ”ML2MT” project from the Volkswagen Stiftung, and the ”The Adaptive Mind” project from the Hessian Ministry of Science and Arts (HMWK). It has further benefited from the HMWK projects ”The Third Wave of Artificial Intelligence - 3AI”, and Hessian.AI, and the research collaboration between TU Darmstadt and Aleph Alpha Research through Lab1141.
Supplementary Materials
Appendix A Experimental Details: Baselines
In this section we provide training details for the three introduced baselines ResNet-18, Pixel-MAC and CBM.
A.1 ResNet-18
For this setup, we initialize the framework with a pretrained ResNet-18 model and employ the Adam optimizer across all experiments. For the CLEVR datasets, we utilize a batch size of 128 and train for 30 epochs with a learning rate of $10^{-4}$ and weight decay of $10^{-4}$. The model is trained using 20 different random seeds, with early stopping based on the validation loss.
For CIFAR-100, we train the ResNet-18 model for 300 epochs using a learning rate of $10^{-4}$ and weight decay of $10^{-4}$, with a batch size of 128. The model is trained across 10 random seeds using early stopping based on validation loss.
For ImageNet, the same setup is applied with a lower learning rate of $10^{-5}$, while keeping the batch size, weight decay, and early stopping criterion unchanged. Both datasets use the standard pixel-level inputs and all trained checkpoints are reused to initialize subsequent Pixel-MAC experiments.
A.2 Pixel-MAC
In this setup, we apply Merlin-Arthur training on pixel space by utilizing the pretrained ResNet-18 models as classifiers and U-Net architectures for both Merlin and Morgana. Throughout all experiments, we employ the Adam optimizer for both classifier and feature selector optimization, with $\gamma=0.5$ to ensure high soundness.
For the CLEVR datasets, we train with a batch size of 128 for 40 epochs, using a learning rate of $10^{-5}$ and weight decay of $10^{-6}$ for the classifier optimization. The U-Net architectures are trained with a learning rate of $10^{-4}$, weight decay of $10^{-5}$ and an L1 penalty coefficient of $0.1$. We set the mask size to 1500, meaning that the U-Nets select a subset of 1500 pixels per sample (out of $128\!\times\!128$ pixels).
For CIFAR-100, we reduce the batch size to 64 and train for 100 epochs. The classifier is optimized with a learning rate of $10^{-5}$ and weight decay of $10^{-4}$. Both Merlin and Morgana are trained using a learning rate of $10^{-3}$ and a reduced mask size of 32 pixels. We use an L1 penalty of $0.01$ and apply early stopping across 10 random seeds.
For ImageNet, we further reduce the batch size to 32 due to memory constraints and train for 80 epochs using the same learning rates and hyperparameters as in the CIFAR-100 setting. The mask size is set to 1000 pixels per image. As with CIFAR-100, early stopping is applied across 10 random seeds, and training is initialized from the pretrained ResNet-18 backbone.
A.3 CBM with Linear Classifier
Next, we present the implementation details for the CBM baseline, where a linear classifier operates on concept features obtained from the concept extractor.
For the CLEVR datasets, we train a linear classifier on concepts extracted by the Neural Concept Binder. The training process employs a batch size of 128, a learning rate of $10^{-3}$, and weight decay of $10^{-4}$. The model is trained for 60 epochs on CLEVR-Hans3 and 30 epochs on CLEVR-Hans7, using early stopping based on validation loss, repeated across 20 different random seeds.
For CIFAR-100 and ImageNet, we train a linear classifier on sparse SpLiCE encodings using a dictionary size of 10,000. In both cases, we use a batch size of 4096 and train for 250 epochs with early stopping, a learning rate of $10^{-3}$, and no weight decay. A hidden layer with 512 units are used, and an L1 penalty of 0.2 is applied within SpLiCE to encourage sparsity in the concept representations. All results are averaged over 10 random seeds.
Appendix B Experiments: NCB-based Neural Concept Verifier
In the following, we provide details on NCB-based NCV, experimental evaluations as well as additional evaluations.
B.1 Pretraining
Before training NCV, we first pretrain the models without the feature selectors. The corresponding results for the pretraining are shown in Table  3, where we evaluate on 20 random seeds. These pretrained models are then used as initialization for the subsequent NCV training. For the pretraining, we use a Set Transformer with two stacked multi-head attention blocks, a hidden dimension of 128 and four attention heads. We use a batchsize of 128, 30 epochs and the Adam optimizer with a learning rate of $10^{-3}$ for both datasets, applying early stopping based on validation loss.
Table 3: Pretraining results on the CLEVR-Hans3 and CLEVR-Hans7 datasets without shortcuts
CLEVR-Hans3 CLEVR-Hans7
Val. Accuracy Test Accuracy Val. Accuracy Test Accuracy
$99.02\pm 0.31$ $98.13\pm 0.37$ $98.08\pm 0.24$ $97.83\pm 0.25$
B.2 NCV Training
For the experiments presented in our main results in Table  1, the experimental details for both datasets are as follows:
Model Architecture. The verifier is implemented as a pretrained Set Transformer consisting of two stacked multi-head attention blocks with hidden dimension 128, four attention heads, and layer normalization. Merlin and Morgana are implemented as independent neural networks, each parameterized by a Set Transformer with two stacked attention blocks with hidden dimensions 256, four attention heads, and layer normalization. The provers receive the full concept slot matrix as input and output a sparse selection mask with exactly 12 nonzero entries (out of 64 total features), indicating the active blocks provided to the verifier.
Training Details. All components are jointly trained using the Adam optimizer with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$. Models are trained for 50 epochs and a batch size of 512 is used throughout. For the Merlin and Morgana provers, a hard selection constraint is enforced, limiting the number of selected concepts to a fixed budget of 12 block-encodings per sample. To ensure high soundness, we set $\gamma=0.5$, giving equal weight to both feature selector losses in the total loss computation. We train our models using 20 random seeds.
Extended Results. Additionally, we evaluated the NCV framework with varying mask sizes and an alternative model architecture for the feature selectors. The results are presented in Table  4 for the CLEVR-Hans3 data set and Table  5 for the CLEVR-Hans7 dataset, where we evaluate both the validation set and the test set. The alternative architecture implements a MLP with two hidden layers and ReLU activation functions for the feature selectors, while maintaining a pretrained Set Transformer as the classifier across all experiments. Our results reveal that the Set Transformer feature selector consistently outperforms the MLP feature selector on test set, particularly with smaller mask sizes such as 4 and 6. Furthermore, this configuration maintains high completeness (>96%) and soundness (>99%), even with a reduced number of selected features.
Table 4: Completeness and soundness on the CLEVR-Hans3 dataset without shortcuts for different mask sizes and feature selector architectures. The highlighted values are used for Table 1.
Validation Test
Mask Size Feature Selector Completeness Soundness Completeness Soundness
4 Set Transformer $98.35\pm 0.31$ $99.88\pm 0.28$ $97.69\pm 0.63$ $99.82\pm 0.23$
MLP $96.51\pm 1.18$ $99.81\pm 0.27$ $95.54\pm 1.37$ $99.85\pm 0.28$
6 Set Transformer $98.71\pm 0.52$ $99.87\pm 0.13$ $98.11\pm 0.62$ $99.88\pm 0.19$
MLP $96.21\pm 0.89$ $99.96\pm 0.03$ $94.78\pm 1.24$ $99.97\pm 0.06$
12 Set Transformer $99.20\pm 0.11$ $100.00\pm 0.00$ $\mathbf{98.92\pm 0.32}$ $\mathbf{100.00\pm 0.00}$
MLP $99.28\pm 0.11$ $99.98\pm 0.06$ $98.89\pm 0.21$ $99.99\pm 0.07$
Table 5: Completeness and soundness on the CLEVR-Hans7 dataset without shortcuts for different mask sizes and feature selector architectures. The highlighted values are used for Table 1.
Validation Test
Mask Size Feature Selector Completeness Soundness Completeness Soundness
4 Set Transformer $96.69\pm 1.28$ $99.93\pm 0.09$ $96.71\pm 1.37$ $99.91\pm 0.09$
MLP $92.63\pm 1.24$ $99.89\pm 0.12$ $92.71\pm 1.31$ $99.87\pm 0.13$
6 Set Transformer $97.32\pm 0.42$ $99.98\pm 0.02$ $97.14\pm 0.51$ $99.98\pm 0.02$
MLP $95.43\pm 1.48$ $99.88\pm 0.13$ $95.12\pm 1.48$ $99.86\pm 0.14$
12 Set Transformer $98.13\pm 0.11$ $100.00\pm 0.00$ $\mathbf{97.89\pm 0.31}$ $\mathbf{100.00\pm 0.00}$
MLP $97.41\pm 1.07$ $99.99\pm 0.03$ $97.01\pm 0.93$ $99.99\pm 0.04$
B.3 Explanations
Here, we present supplementary examples of explanations generated by both Pixel-MAC and NCV on the CLEVR-Hans3 dataset in Figure  4. These results further substantiate our claim that NCV provides significantly more transparent and interpretable explanations compared to the pixel-based PVG baseline.
Figure 4: Comparison of explanations from NCV vs. Pixel-MAC for CLEVR-Hans3 images of all three classes. (a) Merlin–Arthur training on pixel space yields uninformative masks. (b) NCV provides clear explanations by highlighting object features corresponding to the class rule. The single-object images are reconstructions from the respective slots selected by Merlin (prover).
B.4 Shortcut Mitigation
In Table  6, we provide complementary results to those in Table  2, evaluating model robustness on CLEVR-Hans3 under varying degrees of confounding. As in the main experiment, models are trained on datasets with increasing ratios of non-confounded samples, and evaluated on both a confounded validation set and a non-confounded test set.
We observe similar trends: NCV consistently achieves the highest test accuracy across all settings, and its validation–test gap narrows more rapidly than for both the linear and non-linear CBM baselines. Even with minimal clean data (e.g., 5%), NCV achieves strong generalization and a markedly smaller gap. These findings reinforce our conclusions from CLEVR-Hans7 and further highlight NCV’s ability to mitigate shortcut learning through its concept-level PVG formulation.
Table 6: NCV shows increased robustness to shortcuts on CLEVR-Hans3. We report validation accuracy on a set with, and test accuracy on a set without shortcuts, as well as the validation–test gap (lower is better) across models trained with varying amounts of clean data. A smaller gap indicates improved generalization and reduced reliance on shortcuts. NCV consistently achieves the highest test accuracy and among the lowest gaps, confirming its ability to learn robust, faithful decision boundaries under the presence of shortcuts.
Ratio Clean (Samples) Model Val Acc (w/ shortcut) Test Acc (w/o shortcut) Val-Test Gap (↓)
[…]
NCV 99.37± $\mathbf{99.37}\mbox{\scriptsize$\pm\,0.18$}$ 97.11± $\mathbf{97.11}\mbox{\scriptsize$\pm\,0.98$}$ $\mathbf{2.26}$
[…]
NCV 99.59± $\mathbf{99.59}\mbox{\scriptsize$\pm\,0.19$}$ 98.88± $\mathbf{98.88}\mbox{\scriptsize$\pm\,0.37$}$ $\mathbf{0.71}$
[…]
NCV 99.37± $\mathbf{99.37}\mbox{\scriptsize$\pm\,0.28$}$ 98.82± $\mathbf{98.82}\mbox{\scriptsize$\pm\,0.67$}$ $\mathbf{0.55}$
Appendix C Experimental Details: CLIP-based Neural Concept Verifier
In the following section, we present the implementation details of CLIP-based NCV training.
C.1 Pretraining
Once more, before starting with the actual NCV training, we first pretrain the models without the provers (Merlin and Morgana). The corresponding results for the pretraining are shown in Table  7. As textual concept descriptions $T$, we used the top 10,000 most frequent one- and two-word phrases from LAION (Schuhmann et al.,, 2021) captions, following the setup and codebase of Bhalla et al., (2024). For pretraining the verifier, we use a two-layer multilayer perceptron (MLP) with a hidden dimension of 512 and GELU activations (Hendrycks and Gimpel,, 2016) on both CIFAR-100 and ImageNet-1k. For CIFAR-100 and ImageNet, we train the verifier using a batch size of 4096 and a learning rate of $10^{-4}$, applying dropout with a rate of 0.3 and a weight decay of $10^{-4}$. Models are trained for 300 epochs with early stopping (patience 10) based on validation loss. All pretraining is conducted without provers, and the resulting verifiers are used to initialize the CLIP-based NCV training.
Dataset Accuracy (%)
CIFAR-100 85.96
ImageNet-1k 77.07
Table 7: Pretraining accuracy of the verifier (without provers) for CLIP-based NCV on CIFAR-100 and ImageNet-1k.
C.2 NCV Training
For the experiments presented in our main results in Table  1, we detail the training setup separately for CIFAR-100 and ImageNet.
Model Architecture. The verifier (Arthur) is initialized as the pretrained two-layer multilayer perceptron (MLP) described above. Merlin and Morgana are implemented as independent neural networks, each parameterized by a two-layer MLP with hidden dimension 512 and ReLU activations. Both provers receive the full concept activation vector as input and output a sparse selection mask indicating the active concepts that are passed to the verifier.
CIFAR-100. For CIFAR-100, all components are trained jointly for 100 epochs using the Adam optimizer. We set the verifier learning rate to $10^{-4}$, and use $5\times 10^{-4}$ for both Merlin and Morgana. A batch size of 256 is used throughout. Weight decay is set to $0.1$, and a hard mask size of 32 concepts is enforced per input. To incentivise sparse masks, an L1 penalty of 0.1 is applied to the provers. A learning rate scheduler (plateau-based) is employed with a patience of 5, minimum learning rate of $10^{-6}$, and decay factor of 0.001. Early stopping is disabled, and all results are averaged over 10 random seeds.
ImageNet. The ImageNet setup mirrors CIFAR-100 in most aspects. We again train for 100 epochs with a batch size of 256, using the same learning rates for verifier ($10^{-4}$) and provers ($5\times 10^{-4}$), mask size of 32 features, and L1 penalty (0.1). Weight decay is reduced to $0.005$ to improve generalization. The same learning rate scheduler and seed setup are used as in the CIFAR-100 experiments.
Table 8: Ablation results for our method on CIFAR-100 with varying mask sizes. We report mean $\pm$ std over 10 seeds.
Mask Size Completeness Train Completeness Validation Soundness Train Soundness Validation
[…]
Table 9: Ablation results for our method on ImageNet with varying mask sizes. We report mean $\pm$ std over 10 seeds.
Mask Size Completeness Train Completeness Validation Soundness Train Soundness Validation
[…]